Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

pm1weekly_2011-281_2011-287

Mission Operations Days: 2011/281 through 2011/287
October 08, 2011 00:00:00 GMT to October 14, 2011 00:00:00 GMT

Aqua Spacecraft and MODIS Instrument Status:

PM-1 "Aqua" is in Fine Pointing Normal Mode
FM1 MODIS is in B-Side Nominal Science Mode
ComponentState(s)Comment
BlackbodyA Off; B On (285K)Nominal Science Mode
Calibration ElectronicsA Off; B OnNominal Science Mode
Control ProcessorA Off; B OnNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ClosedNominal Science Mode
FDDI FormatterA Off; B OnNominal Science Mode
FDDI PortA Off; B OnNominal Science Mode
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 Off, #2 OnNominal Science Mode
PV VIS FPAA Off; B OnNominal Science Mode
PV NIR FPAA Off; B OnNominal Science Mode
PV SMIR FPAA Off; B OnNominal Science Mode
PV LWIR FPAA Off; B OnNominal Science Mode
PC LWIR FPAA Off; B OnNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOffNominal Science Mode
RC SMIR FPA HeaterOn (83K)Nominal Science Mode
Scan AssemblyA Off; B OnNominal Science Mode
SDSMOffNominal Science Mode
SRCAOff
  • 10W Lamps - 1, 4 (functional)
  • 10W Lamps - 2, 3 (non-functional)
  • 1W Lamps - 1, 2 (functional)
Nominal Science Mode
PS1 Survival HeatersEnabledNominal Science Mode
PS2 Survival HeatersEnabledNominal Science Mode
Timing GeneratorA Off; B OnNominal Science Mode
Flight SoftwareRev BJ + 1 set of patches
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
Nominal Science Mode
Inhibit Ids SetNoneNominal Science Mode

This Week's Completed MODIS Activities:

Saturday, October 08, 2011

None

Sunday, October 09, 2011

None

Monday, October 10, 2011

2011/283 20:06:57-20:20:55 MCL MOD_OA16_SDSM_SCREEN.2
OA-16: SD/SDSM Screened - B-Side
2011/283 21:05:32-21:05:47 MCL MOD_SRCA_SETUP_L1_L4_L5_CC.1
SRCA setup activity using lamps 1, 4, and 5, and in constant current
feedback mode
2011/283 21:08:32-21:38:57 MCL MOD_OA19_SRCA_FULL_RAD.2
OA-19: SRCA Full Radiometric Calibration - B Side
Data collection 21:24:33-21:38:56 (Centered on S/C Night)
S/C Night 21:15:17-21:48:12.

Tuesday, October 11, 2011

None

Wednesday, October 12, 2011

None

Thursday, October 13, 2011

None

Friday, October 14, 2011

None

This Week's Scheduled MODIS Activities Not Completed:

None

This Month's Upcoming MODIS Events:

Calibrations for October will include a Lunar Roll (DOY 280), an SRCA Radiometric (DOY 283), an SDSM Screened (DOY 283), an SDSM Opened (DOY 304), and another SDSM Screened (DOY 304).

Maneuvers:

There will be an Aqua Drag Make Up Maneuver (DMU) #53 on DOY 300 (10/27/2011). Maneuver start time is set for 300/13:55:00 and maneuver end time is set for 300/14:54:10. The maximum yaw angle for the maneuver will reach 14.35 degrees. The projected loss of pointing accuracy times are 300/13:55:00 - 15:09:10.

MODIS Anomalies:

None

General Instrument Comments:

FM1 MODIS is in Science Mode. The SVD and NAD are open, the survival heaters are enabled, and the Black Body is heated to 285 K.

MODIS Telemetry Trends:

The FM1 MODIS instrument telemetry is trending nominally.

Non-MODIS Significant Events:

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.

http://jupiter02.gsfc.nasa.gov/

Limited Life Item Status:

SRCA 10W Lamp #1: 293.9 of 500 hours
SRCA 10W Lamp #2: 188.0 of 500 hours
SRCA 10W Lamp #3: 205.7 of 500 hours
SRCA 10W Lamp #4: 104.8 of 500 hours

SRCA 1W Lamp #1: 523.8 of 5000 hours
SRCA 1W Lamp #2: 274.9 of 5000 hours

Solar Diffuser Door: 2970 of 3022 Movements
Nadir Aperture Door: 1053 of 1316 Movements
Space View Door: 632 of 1316 Movements